# New examples of words for which the binomial complexities and the subword complexity coincide

### v1

@@ Abstract
The complexity of an infinite word can be measured in several ways, the two most common measures being the subword complexity and the abelian complexity. In 2015, Rigo and Salimov introduced a family of intermediate complexities indexed by $k\in\mathbb{N}_{>0}$: the $k$-binomial complexities. These complexities scale up from the abelian complexity, with which the $1$-binomial complexity coincides, to the subword complexity, to which they converge pointwise as $k$ tends to $\infty$. In this article, we provide four classes of $d$-ary infinite words – namely, $d$-ary 11-balanced words, words with subword complexity $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$ (which form a subclass of quasi-Sturmian words), hypercubic billiard words, and words obtained by coloring a Sturmian word with another Sturmian word – for which this scale “collapses”, that is, for which all $k$-binomial complexities, for $k\geq 2$, coincide with the subword complexity. This work generalizes a result of Rigo and Salimov, established in their seminal 2015 paper, which asserts that the $k$-binomial complexity of any Sturmian word coincides with its subword complexity whenever $k\geq 2$.
@@ 1 Introduction
The subword complexity and the abelian complexity are two very classical ways to measure how “complicated” an infinite word is. In 2015, Rigo and Salimov introduced a family of complexities forming a scale between the abelian complexity and the subword complexity: the $k$-binomial complexities [RS15]. These complexities are parameterized by the integer $k\in\mathbb{N}_{>0}$ and are defined via a family of equivalence relations called $k$-binomial equivalences. For $k=1$, this relation corresponds to the abelian equivalence, and the $1$-binomial complexity is the same as the abelian complexity. The $(k+1)$-binomial equivalence is a refinement of the $k$-binomial one, and as $k$ increases, the $k$-binomial equivalence gets progressively closer to equality between finite words. Consequently, the $k$-binomial complexities form a scale from the abelian complexity to the subword complexity, and the $k$-binomial complexity of a given word converges pointwise to its subword complexity as $k\to\infty$.
Since 2015, $k$-binomial complexities have been the subject of several papers [LLR20, LRR20b, LCWW24, RSW24, GRW24]. Interested readers may also consult [RSW23, RRW25] for variations on $k$-binomial complexities and [LRR20a, Whi21] for recent articles on $k$-binomial equivalence relations.
@@
This article contributes to the study of words for which, similarly to Sturmian words, their $k$-binomial complexity coincides with their subword complexity for some small integer $k$. Although we focus mainly on the case $k=2$, several of our results also hold for arbitrary integer $k$ (see Lemmas 8 and 9 below).
@@
In contrast, characterizing words whose $2$-binomial complexity coincides with their subword complexity is considerably more difficult and remains an open problem. In the literature, apart from the Sturmian case and the case of words whose $1$-binomial complexity coincides with their subword complexity, the only other known example is the Tribonacci word [Rau82], which is the most famous and widely studied Arnoux-Rauzy word [AR91].
Proposition 2 ([LRR20b]).
@@
Arnoux-Rauzy words (and, more generally, strict episturmian words [DJP01, GJ09]) form a class of ternary (resp. $d$-ary) infinite words that can be viewed as a combinatorial and arithmetic generalization of Sturmian words. Lejeune, Rigo and Rosenfeld conjectured that all Arnoux-Rauzy words have their $2$-binomial complexity equal to their subword complexity (see [LRR20b] and [Lej21, Chapter 4, Section 4.4 and Appendix B]; see also Section 7 of this article where some numerical experiments are discussed).
@@
The $2$-binomial complexity of any Arnoux-Rauzy word coincides with its subword complexity.
Although the Tribonacci word is the only known example of an Arnoux-Rauzy word satisfying this property, the computer-assisted proof developed for it could potentially be applied to other purely morphic Arnoux-Rauzy words to determine whether this property holds in those cases as well (see [LRR20b, Section 6]).
@@
(iv) If $w$ is obtained as the coloring of a Sturmian word with another Sturmian words [DMP24], then its $2$-binomial complexity is equal to its subword complexity.
Interestingly, all these examples can be thought of as generalizations of Sturmian words. At the time of writing, apart from the trivial case described above where the $1$-binomial complexity coincides with the subword complexity, we are not aware of any word genuinely unrelated to Sturmian words and whose $2$-binomial complexity coincides with its subword complexity.
@@
Finally, let us mention that for every $k\geq 3$, it is known that there exist words whose $k$-binomial complexity, but not their (k1)(k-1)-binomial complexity, coincides with their subword complexity. This was proved by Rigo, Stipulanti and Whiteland in [RSW24], where they showed that the image of any Sturmian word under (k2)(k-2) iterations of the Thue-Morse substitution satisfies this property, see [RSW24, Theorems 7.2 and 7.4].
@@ 2 Definitions and notation
We say that a word $w$ is $c$-balanced ($c\in\mathbb{N}$) if, for every equally long factors $u,v$ of $w$, and for every letter $a\in\mathcal{A}$, we have $\big{|}|u|_{a}-|v|_{a}\big{|}\leq c$. For example, the word $w=11212$ is $1$-balanced while the word $w^{\prime}=11122$ is $2$-balanced but not $1$-balanced (in [And21a, And21b, AV23, AV24], the optimal constant of balancedness is termed imbalance). An infinite word is said to be unbalanced if it is not $c$-balanced for any $c\in\mathbb{N}$.
Given two alphabets $\mathcal{A}$ and $\mathcal{B}$, a substitution $\sigma:\mathcal{A}\to\mathcal{B}^{*}$ is a letters to words application, which we extend to a morphism over $\mathcal{A}^{*}$ (seen as the free monoid for the concatenation). For instance, if $\sigma:\{0,1\}\to\{0,1\}^{*}$ is defined by $\sigma:0\mapsto 01,\,1\mapsto 0$, then $\sigma(010)=\sigma(0)\sigma(1)\sigma(0)=01001$. More generally, substitutions are also extended over the set of infinite words in the following standard way: if $w=a_{0}a_{1}a_{2}a_{3}\ldots$ where each $a_{i}$ is a letter from the alphabet $\mathcal{A}$, then $\sigma(w):=\sigma(a_{0})\sigma(a_{1})\sigma(a_{2})\sigma(a_{3})\ldots$.
@@
The subword complexity (resp. abelian complexity and $k$-binomial complexity) of a word $w$ is the function $p_{w}:\mathbb{N}\to\mathbb{N}$ (resp. $\rho_{w}:\mathbb{N}\to\mathbb{N}$ and $b_{w}^{k}:\mathbb{N}\to\mathbb{N}$) that counts, for each integer $n\in\mathbb{N}$, the number of distinct factors (resp. abelian classes and $k$-binomial classes) of $w$ of length111 Since two factors are abelian (resp. $k$-binomially) equivalent only if they are of the same length, each abelian (resp. $k$-binomial) class contains only equally long factors. The common length of the factors of a given abelian (resp. $k$-binomial) class is referred to as the length of the class. $n$:
$\begin{array}[]{rcclrcclrccl}p_{w}:&\mathbb{N}&\rightarrow&\mathbb{N}&\rho_{w}:&\mathbb{N}&\rightarrow&\mathbb{N}&b_{w}^{k}:&\mathbb{N}&\rightarrow&\mathbb{N}\\ &n&\mapsto&\#\big{(}\mathcal{L}_{n}(w)\big{)}&&n&\mapsto&\#\big{(}\mathcal{L}_{n}(w)/_{\sim_{\text{ab}}}\big{)}&&n&\mapsto&\#\big{(}\mathcal{L}_{n}(w)/_{\sim_{k}}\big{)}\end{array}$
@@
$w_{\mathtt{fibo}}=01001010010010100101001001010010010100100101001010\ldots$
@@
A consequence of Morse and Hedlund’s Theorem – which states that an infinite word $w$ is ultimately periodic if and only if there exists $m\in\mathbb{N}$ such that $p_{w}(m)\leq m$ [MH38] – is that Sturmian words are exactly the (binary) non-eventually periodic words with minimal complexity. Moreover, the minimal subword complexity of a $d$-ary non-eventually periodic word is given by $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$. Ferenczi and Mauduit provided a characterization of such words in [FM97, Lemma 1 and Lemma 4].
More generally, an infinite word $w$ is called quasi-Sturmian if there exist two integers $n_{0},k_{0}$ such that $p_{w}(n)=n+k_{0}$ for every $n\geq n_{0}$. These words have been characterized by Cassaigne in [Cas97, Proposition 8].
@@
Coloring of words. Given two infinite words $w_{0}\in\mathcal{A}^{\mathbb{N}}$ and $w_{1}\in\mathcal{B}^{\mathbb{N}}$ written over disjoint alphabets ($\mathcal{A}\cap\mathcal{B}=\emptyset$), we define, for every letter $a\in\mathcal{A}$, a new infinite word $\mathtt{color}(w_{0},a,w_{1})\in(\mathcal{A}\sqcup\mathcal{B}\setminus\{a\})^{\mathbb{N}}$ obtained as the coloring of the letter $a$ in $w_{0}$ by $w_{1}$. Formally, we set
@@
$\begin{array}[]{rcl}w_{0}&=&\mathtt{0a00a0a00a00a0a00a0a00a00a0a00a00a0a00a00a0a00a0a0}\ldots\\ w_{1}&=&\mathtt{{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}}\ldots\\ \mathtt{color}(w_{0},\mathtt{a},w_{1})&=&\mathtt{01002010010020100201001002010010020100200101002010}\ldots\end{array}$
@@
Coloring of words can be used to characterize non-eventually periodic $1$-balanced words; indeed, any such word can be obtained as the coloring of a Sturmian word by one or two words with constant gaps [Gra73, Hub00]. More recently, Dvořáková, Masáková and Pelantová introduced a new class of ternary words obtained as the coloring of a Sturmian word by another Sturmian word [DMP24]. These words share several properties with cubic billiard words, such as quadratic complexity [AMST94a, AMST94b]222 There is an error in these articles, see [Béd03] for a correction. Interested readers may also consult [Bar95, Béd09] for a generalization of this result to arbitrary dimension $d$. , $2$-balancedness [Vui03], and an eventually constant abelian complexity, equal to $4$ [AV23]. Nonetheless, these two classes of words do not coincide. This can be seen in several ways.
@@
    Subword complexity: A cubic billiard word with rationally independent letter frequencies, whose inverses are also rationally independent, has subword complexity $n\mapsto n^{2}+n+1$, [Bar95] (see also [Béd03, Theorem 6]). In contrast, the subword complexity of the coloring of a Sturmian word by another Sturmian word is bounded by $\alpha n^{2}(1+o(1))$, where $\alpha\in(0,1)$ is the frequency of the colored letter in the original Sturmian word [DMP24, Theorem 9 and Remark 10].
    Abelian complexity: The abelian complexity of a cubic billiard word with rationally independent letter frequencies is equal to $4$ for every length $n\geq 2$ [AV23, Theorem 8]. On the other hand, if the coloring of a Sturmian word by another Sturmian word has rationally independent letter frequencies, then its abelian complexity only reaches $4$ when $n\geq\lceil\alpha^{-1}\rceil$, where $\alpha$ stills denote the frequency of the colored letter in the original Sturmian word [DMP24, Theorem 16 and its proof].
    Binary projections: All the binary projections of a cubic billiard words are $1$-balanced, but this is not generally true for the coloring of a Sturmian word by another Sturmian word. For example, the coloring of the Fibonacci word by the Fibonacci word above yields the projection,
    $\pi_{\mathtt{0},\mathtt{1}}(\mathtt{color}(w_{0},\mathtt{a},w_{1}))=\mathtt{0100010010001000100100010010001000010100010}\ldots$
    which is not $1$-balanced, as it contains the factors $\mathtt{000}$ and $\mathtt{101}$.

### v2

# New examples of words for which binomial complexities and subword complexity coincide

@@ Abstract
The complexity of an infinite word can be measured in several ways, the two most common measures being the subword complexity and the abelian complexity. In 2015, Rigo and Salimov introduced a family of intermediate complexities indexed by $k\in\mathbb{N}_{>0}$: the $k$-binomial complexities. These complexities scale up from the abelian complexity, with which the $1$-binomial complexity coincides, to the subword complexity, to which they converge pointwise as $k$ tends to $\infty$. In this article, we provide four classes of $d$-ary infinite words—namely, $d$-ary 11-balanced words, words with subword complexity $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$ (which form a subclass of the so-called quasi-Sturmian words), hypercubic billiard words, and words constructed by repeated Sturmian colorings—for which this scale “collapses”, that is, all $k$-binomial complexities, for $k\geq 2$, coincide with the subword complexity. This work generalizes a result of Rigo and Salimov, established in their seminal paper from 2015, which asserts that the $k$-binomial complexity of any Sturmian word coincides with its subword complexity whenever $k\geq 2$.
@@ 1 Introduction
The subword complexity and the abelian complexity are two very classical ways to measure how “complicated” an infinite word is. In 2015, Rigo and Salimov introduced a family of complexities forming a scale between the abelian complexity and the subword complexity: the $k$-binomial complexities [RS15]. These complexities are parameterized by the integer $k\in\mathbb{N}_{>0}$ and are defined via a family of equivalence relations called $k$-binomial equivalences. For $k=1$, this relation corresponds to the abelian equivalence, and the $1$-binomial complexity is the same as the abelian complexity. The $(k+1)$-binomial equivalence is a refinement of the $k$-binomial one, and as $k$ increases, the $k$-binomial equivalence gets progressively closer to equality between finite words. Consequently, the $k$-binomial complexities form a scale from the abelian complexity to the subword complexity, and the $k$-binomial complexity of a given word converges pointwise to its subword complexity as $k$ tends to $\infty$.
Since 2015, $k$-binomial complexities have been the subject of several papers [LLR20a, LRR20b, LCW+24, RSW24, GRW24]. Interested readers may also consult [RSWip, RRW25] for variations on $k$-binomial complexities and [LRR20c, WHI21] for recent articles on $k$-binomial equivalence relations.
@@
This article contributes to the study of words for which, similarly to Sturmian words, the $k$-binomial complexity coincides with the subword complexity for some small integer $k$. Although we focus mainly on the case $k=2$, several of our results also hold for arbitrary integer $k$ (see Lemmas 8 and 9 below).
@@
By contrast, characterizing words whose $2$-binomial complexity coincides with their subword complexity is considerably more difficult and remains an open problem. In the literature, apart from the Sturmian case and the case of words whose $1$-binomial complexity coincides with their subword complexity, the only other known example is the Tribonacci word [RAU82], which is the most famous and widely studied instance of the class of Arnoux–Rauzy words [AR91].
Proposition 2 ([LRR20b] Theorem 29).
@@
Arnoux–Rauzy words (and, more generally, strict episturmian words [DJP01, GJ09]) form a class of ternary (resp. $d$-ary) infinite words that can be viewed as a combinatorial and arithmetic generalization of Sturmian words. Lejeune, Rigo and Rosenfeld conjectured that all Arnoux–Rauzy words have their $2$-binomial complexity equal to their subword complexity (see [LRR20b] and [LEJ21, Chapter 4, Section 4.4 and Appendix B]; see also Section 7 of this article where some numerical experiments are discussed).
@@
The $2$-binomial complexity of any Arnoux–Rauzy word coincides with its subword complexity.
Although the Tribonacci word is the only known example of an Arnoux–Rauzy word satisfying this property, the computer-assisted proof developed for it could potentially be applied to other purely morphic Arnoux–Rauzy words to determine whether this property holds in those cases as well (see [LRR20b, Section 6]).
@@
(iv) If $w$ is a $d$-ary Sturmian coloring, then its $2$-binomial complexity is equal to its subword complexity.
These four classes of words are defined in Section 2. Interestingly, all these examples can be thought of as generalizations of Sturmian words. At the time of writing, apart from the trivial case described above where the $1$-binomial complexity coincides with the subword complexity, we are not aware of any word genuinely unrelated to Sturmian words and whose $2$-binomial complexity coincides with its subword complexity.
@@
Finally, let us mention that for every $k\geq 3$, it is known that there exist words whose $k$-binomial complexity, but not their (k1)(k-1)-binomial complexity, coincides with their subword complexity. This was proved by Rigo, Stipulanti and Whiteland in [RSW24], where they showed that the image of any Sturmian word under (k2)(k-2) iterations of the Thue–Morse substitution satisfies this property, see [RSW24, Theorems 7.2 and 7.4].
@@ 2.1 Basics and complexities
We say that a word $w$ is $c$-balanced ($c\in\mathbb{N}$) if, for every equally long factors $u,v$ of $w$, and for every letter $a\in\mathcal{A}$, we have $\big||u|_{a}-|v|_{a}\big|\leq c$. For example, the word $w=11212$ is $1$-balanced while the word $w^{\prime}=11122$ is $2$-balanced but not $1$-balanced. The optimal constant of balancedness is termed imbalance in [AND21b, AND21a, AV22, AV23, AV24]. An infinite word is said to be unbalanced if it is not $c$-balanced for any $c\in\mathbb{N}$.
Given two alphabets $\mathcal{A}$ and $\mathcal{B}$, a substitution $\sigma:\mathcal{A}\to\mathcal{B}^{*}$ is a letter-to-word application, which we extend to a morphism over $\mathcal{A}^{*}$ (seen as the free monoid for the concatenation). For instance, if $\sigma:\{0,1\}\to\{0,1\}^{*}$ is defined by $\sigma:0\mapsto 01,\,1\mapsto 0$, then $\sigma(010)=\sigma(0)\sigma(1)\sigma(0)=01001$. More generally, substitutions are also extended over the set of infinite words in the following standard way: if $w=a_{0}a_{1}a_{2}a_{3}\ldots$ where each $a_{i}$ is a letter from the alphabet $\mathcal{A}$, then $\sigma(w):=\sigma(a_{0})\sigma(a_{1})\sigma(a_{2})\sigma(a_{3})\ldots$.
@@
The subword complexity (resp. abelian complexity and $k$-binomial complexity) of a word $w$ is the function $p_{w}:\mathbb{N}\to\mathbb{N}$ (resp. $\rho_{w}^{ab}:\mathbb{N}\to\mathbb{N}$ and $b_{w}^{k}:\mathbb{N}\to\mathbb{N}$) that counts, for each integer $n\in\mathbb{N}$, the number of distinct factors (resp. abelian classes and $k$-binomial classes) of $w$ of length [^1] Since two factors are abelian (resp. $k$-binomially) equivalent only if they are of the same length, each abelian (resp. $k$-binomial) class contains only equally long factors. The common length of the factors of a given abelian (resp. $k$-binomial) class is referred to as the length of the class. $n$:
$\begin{array}[]{rcclrcclrccl}p_{w}:&\mathbb{N}&\rightarrow&\mathbb{N}&\rho_{w}^{ab}:&\mathbb{N}&\rightarrow&\mathbb{N}&b_{w}^{k}:&\mathbb{N}&\rightarrow&\mathbb{N}\\ &n&\mapsto&\#\big(\mathcal{L}_{n}(w)\big)&&n&\mapsto&\#\big(\mathcal{L}_{n}(w)/_{\sim_{\text{ab}}}\big)&&n&\mapsto&\#\big(\mathcal{L}_{n}(w)/_{\sim_{k}}\big)\end{array}$
@@ 2.2 Classes of words
$w_{\mathtt{fibo}}=01001010010010100101001001010010010100101001001010\ldots$
@@
A consequence of Morse and Hedlund’s Theorem—which states that an infinite word $w$ is eventually periodic if and only if there exists $m\in\mathbb{N}$ such that $p_{w}(m)\leq m$ [MH38]—is that Sturmian words are exactly the (binary) non-eventually periodic words with minimal complexity. Moreover, the minimal subword complexity of a $d$-ary non-eventually periodic word is given by $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$ (by $d$-ary, we require that $d$ distinct letters appear in the word). Ferenczi and Mauduit provided a characterization of such words in [FM97, Lemma 1 and Lemma 4].
More generally, an infinite word $w$ is called quasi-Sturmian if there exist two integers $n_{0},k_{0}$ such that $p_{w}(n)=n+k_{0}$ for every $n\geq n_{0}$. These words have been characterized by Cassaigne in [CASd, Proposition 8].
@@
Coloring of words. Given two infinite words $w_{0}\in\mathcal{A}^{\mathbb{N}}$ and $w_{1}\in\mathcal{B}^{\mathbb{N}}$ written over disjoint alphabets ($\mathcal{A}\cap\mathcal{B}=\emptyset$), we define, for every letter $a\in\mathcal{A}$, a new infinite word $\mathtt{color}(w_{0},a,w_{1})\in(\mathcal{A}\sqcup\mathcal{B}\setminus\{a\})^{\mathbb{N}}$ obtained by coloring the letter $a$ in $w_{0}$ with $w_{1}$. Formally, we set
@@
$\begin{array}[]{rcl}w_{0}&=&\mathtt{0a00a0a00a00a0a00a0a00a00a0a00a00a0a00a0a00a00a0a0}\ldots\\ w_{1}&=&\mathtt{{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}{\color[rgb]{1,1,1}\mathtt{0}}2{\color[rgb]{1,1,1}\mathtt{0}}1{\color[rgb]{1,1,1}\mathtt{0}}}\ldots\\ \mathtt{color}(w_{0},\mathtt{a},w_{1})&=&\mathtt{01002010010020100201001002010010020100201001002010}\ldots\end{array}$
@@
In the literature, colorings of words are used, for example, to characterize non-eventually periodic $1$-balanced words. Indeed, any such word can be obtained as the coloring of a Sturmian word by one or two words with constant gaps [GRA73, HUB00]. More recently, Dvořáková, Masáková, Mašek, and Pelantová introduced a new class of ternary and quaternary words obtained by coloring a Sturmian word with one or two Sturmian words [DMP23, DMP24]. The word $\mathtt{color}(w_{0},\mathtt{a},w_{1})$ defined above belongs to this class. These words share several properties with cubic billiard words, such as quadratic complexity [AMS+94a, AMS+94b, BÉD03] (interested readers may also consult [BAR95, BÉD09] for a generalization of this result to arbitrary dimension $d$), 22-balancedness [VUI03], and an eventually constant abelian complexity, equal to $4$ [AV23].
More generally, we define $d$-ary Sturmian colorings iteratively as follows:
    11-ary Sturmian colorings are the infinite constant words (e.g., $0000000\ldots$),
    a $d$-ary word is a Sturmian coloring if it can be written as $\mathtt{color}(w_{0},a,w_{1})$, where $w_{0}\in\mathcal{A}^{\mathbb{N}}$ is a (d1)(d-1)-ary Sturmian coloring, $a\in\mathcal{A}$, and $w_{1}$ is a Sturmian word over a binary alphabet $\mathcal{B}$ disjoint from $\mathcal{A}$ (i.e., $\mathcal{A}\cap\mathcal{B}=\emptyset$).
It follows from this definition that binary Sturmian colorings are exactly Sturmian words.
2.3 Comparison between the four classes of words in Theorem 1.
The four classes of words this article focuses on are widely studied in the literature and are presented as generalizations of Sturmian words. Note that none of these classes is contained in another. Indeed, if we set
[…]
then:
    $C_{1}\nsubseteq C_{2}\cup C_{3}\cup C_{4}$. Indeed, $w:=3(12)^{\omega}$ belongs to $C_{1}$ but not to $C_{2}$ (its subword complexity is bounded), nor to $C_{3}$ or $C_{4}$ since it is not recurrent.
    $C_{2}\nsubseteq C_{1}\cup C_{3}\cup C_{4}$. Let $w:=3\cdot S(w_{0})=3211212112112\ldots$, where $w_{0}\in\{1,2\}^{\mathbb{N}}$ is the Fibonacci word and $S$ denotes the shift operator, i.e., the operator that erases the first letter of an infinite word. Then $w$ belongs to $C_{2}$ but not to $C_{1}$ (it contains both factors $32$ and $11$), nor to $C_{3}$ or $C_{4}$ since it is not recurrent.
@@
    $C_{3}\nsubseteq C_{1}\cup C_{2}\cup C_{4}$. Let $w$ be a cubic billiard word with rationally independent letter frequencies whose inverses are also rationally independent (note that this is the generic situation). Then its subword complexity is given by $n\mapsto n^{2}+n+1$ [BAR95] (see also [BÉD03, Theorem 6]), which prevents it from belonging to $C_{2}$ or $C_{4}$ (the subword complexity of a ternary Sturmian coloring is bounded by $\alpha n^{2}(1+o(1))$, where $\alpha\in(0,1)$ denotes the frequency of the colored letter in the original Sturmian word [DMP24, Theorem 9 and Remark 10]). Furthermore, $w$ is not $1$-balanced [AV22, Theorem 2].
    $C_{4}\nsubseteq C_{1}\cup C_{2}\cup C_{3}$. Let $w$ be a ternary Sturmian coloring with rationally independent letter frequencies whose inverses are also rationally independent (this is again the generic situation). Then its subword complexity satisfies $p_{w}(n)\sim_{n\to\infty}\alpha n^{2}$, where $\alpha$ again denotes the frequency of the colored letter in the original Sturmian word [DMP24, Theorem 16 and Remark 10], which prevents it from belonging to $C_{2}$ or $C_{3}$. Moreover, $w$ is not $1$-balanced [DMP24, Theorem 9].
However, these classes are not disjoint. An interesting example is the following. Let $w_{0}\in\{0,1\}^{\mathbb{N}}$ be the Fibonacci word (or any other Sturmian word), and set $w_{1}:=(23)^{\omega}$. One can show that the colored word $w:=\mathtt{color}(w_{0},0,w_{1})$ belongs to $C_{1}\cap C_{3}\cap C_{4}$. In other words, $w$ is simultaneously a $1$-balanced ternary word, a cubic billiard word, and the coloring of a Sturmian word by another Sturmian word (of course, for a coloring different from the one defining $w$).